Problem 23-4A (Part Level Submission) Kansas Company uses a standard cost accounting system. In 2017, the company produced 27,600 units. Each unit took several pounds of direct materials and 1.6 standard hours of direct labor at a standard hourly rate of $13.00. Normal capacity was 49,700 direct labor hours. During the year, 130,800 pounds of raw materials were purchased at $0.91 per pound. All materials purchased were used during the year. (a) Your answer is correct. If the materials price variance was $5,232 favorable, what was the standard materials price per pound? (Round answer to 2 decimal places, e.g. 2.75.) Standard materials price per pound $ Click if you would like to Show Work for this question: Open Show Work Show Solution Show Answer Link to Text Link to Text Attempts: 1 of 3 used (b) Your answer is correct. If the materials quantity variance was $14,136 unfavorable, what was the standard materials quantity per unit? (Round answer to 1 decimal place, e.g. 1.5.) Standard materials quantity per unit Click if you would like to Show Work for this question: Open Show Work Show Solution Show Answer Link to Text Link to Text Attempts: 2 of 3 used (c) Your answer is correct. What were the standard hours allowed for the units produced? Standard hours allowed Click if you would like to Show Work for this question: Open Show Work Show Solution Show Answer Link to Text Link to Text Attempts: 1 of 3 used (d) Your answer is correct. If the labor quantity variance was $5,200 unfavorable, what were the actual direct labor hours worked? Actual hours worked Click if you would like to Show Work for this question: Open Show Work Show Solution Show Answer Link to Text Link to Text Attempts: 1 of 3 used (e) Your answer is correct. If the labor price variance was $8,912 favorable, what was the actual rate per hour? (Round answer to 2 decimal places, e.g. 2.75.) Actual rate per hour $ Click if you would like to Show Work for this question: Open Show Work Show Solution Show Answer Link to Text Link to Text Attempts: 1 of 3 used (f) Your answer is correct. If total budgeted manufacturing overhead was $323,050 at normal capacity, what was the predetermined overhead rate? (Round answer to 2 decimal places, e.g. 2.75.) Predetermined overhead rate $ Click if you would like to Show Work for this question: Open Show Work Show Solution Show Answer Link to Text Link to Text Attempts: 1 of 3 used (g) What was the standard cost per unit of product? (Round answer to 2 decimal places, e.g. 2.75.) Standard cost per unit $
Answer #1
1) If the materials price variance was $5,232 favorable, what was the standard materials price per pound? (Round answer to 2 decimal places, e.g. 2.75.)
(Actual price - Standard price) x Actual quantity used = Material price variance
($0.91 - SP ) x 130800 pounds = - $5232
(130800 x $.91 +$5232 )/130800
Standard materials price per pound $           0.95
2) If the materials quantity variance was $14,136 unfavorable, what was the standard materials quantity per unit? (Round answer to 1 decimal place, e.g. 1.5.)
Standard Usage in Units = $14136/$.95)      14,880.00 Pounds
Standard Quantity for 27600 units = (130,800 - 14880)pounds 1,15,920.00 Pounds
Standard materials quantity per unit = (115920/27600 units)               4.20 Pounds per unit
3) What were the standard hours allowed for the units produced?
Standard hours allowed = standard hours of direct labor x units produced
Standard hours allowed = 1.6 x 27,600 units      44,160.00 Hours
4) If the labor quantity variance was $5,200 unfavorable, what were the actual direct labor hours worked?
DL Efficiency Variance = ( SH − AH ) × SR
-$5200 = (44,160 - AH) x $13
Actual hours worked = (44160 x $13 + 5200)$13      44,560.00 Hours
5) If the labor price variance was $8,912 favorable, what was the actual rate per hour? (Round answer to 2 decimal places, e.g. 2.75.)
(Actual rate - Standard rate) x Actual hours worked = Labor rate variance
(AR - $13) x 44560 = $8912
Actual rate per hour = (($13 x 44560)-8912)/44560 $         12.80
6)
If total budgeted manufacturing overhead was $323,050 at normal capacity, what was the predetermined overhead rate?
Predetermined overhead rate = 323050/49700 DL 6.5 Per direct labour hour
7)
What was the standard cost per unit of product?
Direct Material = 4.20 pound per unit x $.95 3.99
Direct Labour = 1.6 x $13 20.8
Manufacturing Overhead = 1.6 x $6.5 10.4
Standard cost per unit $         35.19